# Mind the Prompt: A Novel Benchmark for Prompt-based Class-Agnostic Counting

### v1

@@ Appendix D TFPOC Density Maps Creation
TFPOC is a training-free method that detects objects to count using the powerful SAM model [13]. Therefore, it is a detection-based model that counts the number of output detections. For this reason, it never really computes a density map, which is the main output interface used to prepare the predictions for the mosaic test and produce the qualitative visualization. To prepare the density maps, we simply plotted the region centers as small dots, each having an area of 1 (as is usually done for preparing ground truth density maps from dot annotations).
@@ Appendix E More Qualitative Results
In Figure 6, we report four images that we used to feed the model together with different negative classes. We can notice how all the methods except DAVE actually count the negative classes, predicting a number of instances comparable to – and in many cases even higher than – the ground truth of the positive class. Notice how DAVE identically predicts zero instances thanks to the proposed inference modification.
In Figure 7, we report more outcomes from the mosaic test. These further examples show how the models struggle to count only the correct class. Notably, despite DAVE being very good at distinguishing the only positive class from the negative test and achieving remarkable results on the PrACo metrics on the mosaic test, it still seems to sometimes catastrophically fail on the mosaic test, switching the correct class with the negative one.

### v2

@@ Appendix D TFPOC Density Maps Creation
TFPOC is a detection-based method that localizes objects to count using the powerful SAM model [16]. For this reason, it never really computes a density map, which is the main output interface used to prepare the predictions for the mosaic test and produce the qualitative visualization. To prepare the density maps, we simply plotted the region centers as small dots, each having an area of 1 (as is usually done for preparing ground truth density maps from dot annotations).
@@ Appendix E More Qualitative Results
In Fig. 6, we present four images provided as input to the model, each paired with different negative classes. Notably, all methods except DAVE count the negative classes, often predicting a number of instances comparable to – or even exceeding – the ground truth for the positive class. In contrast, DAVE consistently predicts zero instances, demonstrating the effectiveness of the proposed inference modification.
In Fig. 7, we present additional results from the mosaic test, illustrating how the models often struggle to count exclusively the correct class. Notably, while DAVE demonstrates strong performance in distinguishing the sole positive class from negative ones and achieves impressive results on the PrACo metrics for the mosaic test, it occasionally suffers catastrophic failures, incorrectly swapping the positive class with a negative one.